OpenStudy (anonymous):

the roots of the equation 3x^2- 27x=0 are

OpenStudy (campbell_st):

take out 3x as a common factor 3x(x - 9) = 0 then x = 0, 9

OpenStudy (callisto):

Factorise the left side, 3x^2 - 27x =0 3x (x-9)=0 So, you can get 3x =0 or x-9 =0 Time to work it out yourself :)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):